Database last updated on Fri, 06 Feb 2026 23:10:51 +0000 by localhost

Critical warnings from b-files' comments for days between 2022081 and 2022081 (1 sec, 5 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
menu:Brewer failed to respond 5 times21B08122.187TORON#18714:28:291-1
re:Reset failed 11 times. Quitting and going to the menu.11B08122.187TORON#18715:16:221-1
hp:Spectrometer alignment trouble.21B08122.187TORON#18719:33:411-1
hg:Lamp not on... hg test terminated F1= 1 F4= 011B08122.187TORON#18719:36:301-1
hg:Lamp not detected f1= 1 f4= 011B08122.187TORON#18719:36:311-1

Instrument constant changes for the week before 2022081 (1 sec, 0 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
187202208133.5

Non-critical warnings from b-files' comments for days between 2022081 and 2022081 (0 sec, 3 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
re:Azimuth discrepancy = -8023299B08122.187TORON#18715:35:56-8023-3030
hp:HOME key pressed199B08122.187TORON#18719:38:281-1
hp:Going to menu199B08122.187TORON#18719:38:301-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhghprssl
081TORONTO#187menu,Must do HG now,Starting schedule inside,Waiting until ZA = 170 now 125.7,Waiting until ZA = 170 now 126.9,Waiting until ZA = 170 now 128,Waiting until ZA = 170 now 129.1,Waiting until ZA = 170 now 130.2,Waiting until ZA = 170 now 131.1,Waiting until ZA = 170 now 131.9,Waiting until ZA = 170 now 132.7,Waiting until ZA = 170 now 133.4,Waiting until ZA = 170 now 133.9,Waiting until ZA = 170 now 134.4,Waiting until ZA = 170 now 134.7,Waiting until ZA = 170 now 135,Waiting until ZA = 170 now 135.152171846
(0 sec)

Relative and absolute humidity for day 2022081 (1 sec, 1 rows)

SerialRH, %AH, g/m37-day RH historyDes-nt replcd days ago
18717.24.0